OpenStudy (anonymous):

I need help The table shows the outputs y for different inputs x: Input (x) 2 4 6 8 Output (y) 1 2 3 4 Part A: Do the data in this table represent a function? Justify your answer. Part B: Compare the data in the table with the relation f(x) = 3x – 10. Which relation has a greater value when x = 8? (2 points) Part C: Using the relation in Part B, what is the value of x if f(x) = 80? What do I do? Medal will be given

OpenStudy (ghostgate):

Well, we have these variables in the 'input' or 'x' and 'output' or 'y' sections: x = 2, 4, 6, 8 y = 1, 2, 3, 4 What is one thing these sets have in common in difference? By difference you have: (This is what I call them not what they actually are known as) Addition Modifier - Where x has 3, 5, 7 and y has 5, 7, 9 - The modifier is +2. Subtraction Modifier - Where x has 5, 7, 9 and y has 3, 5, 7 - The modifier is -2. Multiplication Modifier - Where x has 3, 5, 7 and y has 6, 10, 14 - The modifier is *2. Division Modifier - Where x has 6, 10, 14 and y has 3, 5, 7 - The modifier is /2.

OpenStudy (anonymous):

okay so I know what the answer to a is, but i am confused with how to solve b

OpenStudy (anonymous):

and c

OpenStudy (ghostgate):

So you figured out that A is /2, since each x variable divided by 2 is the y variables. And you figured out that C is 40, since you know that the difference is /2, right?

OpenStudy (anonymous):

still confused... sorry

OpenStudy (ghostgate):

Well you said you figured out Part A and Part C, right? I was just going over those answers to make sure we were on the same page. As for Part B: Hmmm, if the question is: Compare the data in the table with the relation f(x) = 3x – 10. Which relation has a greater value when x = 8? Well, if I guess by relations it refers to: f(x) = 3x - 10 and f(x) = x/2 I think. So... if this is the case then I would say that f(x) = 3x - 10 is the answer since: f(x) = 3(9) - 10 -> 27 - 10 = 17 f(x) = 9/2 = 4.5 However, let's do a larger number: f(x) = 3(35) - 10 -> 105 - 10 = 95 f(x) = 35/2 = 17.5 Well, I'm not sure if that's the answer. I hope this helped in someway! Have a great day! {---Ghostgate---}

OpenStudy (anonymous):

Thank you so much, really appreciated :)

OpenStudy (ghostgate):

You're welcome! Hopefully it's the answer, if you need any help in the future, tag me, I'll help as best as I can, I'm only really good at stuff like Math and Writing though. ;)